OpenStudy (anonymous):

the length of a rectangle is 1cm more than the widtrh. if the perimeter of the rectangle is 6cm. what are the dimension of the rectangle.?

OpenStudy (anonymous):

another linear equations problem y = x + 1 2(y + x) = 6 substitute in y, solve. :)

OpenStudy (anonymous):

y = x + 1 2(y + x) = 6 2(2x+1)=6 2x+1=3 2x=2 x=1 y=x+1=2